OpenStudy (anonymous):

suppose the function below was expressed in standard form Y=ax^2+bx+c. what is the value of a with the vertex -3,-2

OpenStudy (amistre64):

if we were to complete the square: y = ax^2 + bx + c y/a = x^2 + b/a x + c/a y/a = (x^2 + b/a x +b^2/4a^2) - b^2/4a^2 + 4ac/4a^2 y/a = (x - (-b/2a))^2 - (b^2-4ac)/4a^2 y = a(x - (-b/2a))^2 - (b^2-4ac)/4a

OpenStudy (anonymous):

thanks.!

OpenStudy (amistre64):

youre welcome in this format we can relate the vertex (p,q) as: y = a(x-p)^2 + q which should give us a means of working out the finer details